Run these commands in R, then describe what the resulting numbers represent.

a) pbinom(q=5, size=10, prob=1/6)

b)

n=10

p=.5

x=9

pbinom(x, n, p)

c) punif(5, min=1, max=10) - punif(4, min=1, max=10)
Answer #1

TOPIC:R-codes and description.

2 pbinom (q=5, size = 10, prob = 15) = 902. = To. 997.5618). Description :- The above command indicates - P (X<5). - where ,so, x = required x-value = 9, = No. of trials = 10, p = p (of success) = 0.5. 3 bunij (5, min= 1, max 10)-bunif (4, minal, ma

REQUIRED R-CODES AND OUTPUTS:

pbinom (q=5, size=10, prob=1/6) [1] 0.9975618 n=10 > pr.5 > x=9 > pbinom (X,n,p) [1] 0.9990234 min=1; max=10) > punif (5, min